WAC Notes July 17 2019

From PREX Wiki
Jump to: navigation, search
Documentation
HOW TOs for shift crew
Expert Tools
All Expert Contacts



PREX Main << Weekly Analysis Coordinator << Instructions

No Prior Day << >> WAC Notes July 18 2019

July 17 2019 Owl Shift

  • WAC = Cameron Clarke
  • Slug 3 (refer to these runs as slug 3.1)
  • Beam Status
    • OWL had ~7 hours of beam, with some sketchy BPM noise and clear BCM DD problems
  • Major issue: BCM Upstairs Black box noise on AN_US

Production Run List (status at 6pm 7/17/2019)

Slug 3

  • Cut first 30k events out: 3398
  • Good: 3399
  • Good: 3400
  • Good: 3401
  • Good: 3402
  • Needs cuts (some miniruns have large RMS): 3403

Slug 4 (part 1)

  • Needs cuts: Beam excursion, (fixed) minirun 6 has absurd RMS, (fixed) BPM noise: 3404
  • Good: 3405
  • Good: 3406
  • Good: 3407

Slug 4 (part 2)

Non-Production Runs and Bad Runs

Pedestal/Calibrations Updates

  • Tao's Pedestal Update
    • Using: prexCH_beamline_pedestal.3396.map for 50uA runs from slug 3 onward - new BPMs in ARC need attention
    • Using: prexCH_beamline_pedestal.3395.map - use it as a template pedestal map for 70 uA runs.
    • I've taken prexCH_beamline_pedestal.3396.map and placed it in the prexCH_beamline_pedestal.3390-.map
  • Caryn's Hall C and cav4cQ pedestals (3373 and 3378 respectively)

Issues

  • Tuning 3397 - HALOG
  • Charge Feedback glitch: non-fatal QW Feedback process fighting with Analysis Minder - HALOG, HALOG
  • OPS FFB on: for a few minutes of run 3398 - HALOG
  • Strange beam motion at beginning of 3398 - HALOG
  • BCM Noise/phase lag question: Begins run 3400, varies, investigated on scope, looks like BCM black box issue - HALOG, Sakib noted, Victoria noted
    • Looked at on a scope with Tune Beam and changing all cables isolates it to a electronics/hardware issue in the Upstairs black box, not downstairs.
    • This will need to be fixed, but probably the data is fine.
  • BPMs have strange noise, especially 4eY: Beings run 3399 - HALOG
  • Compliment BPMs don't have this noise (anecdotal), and their DDs w.r.t. normal BPMs is noticeable - HALOG, HALOG 2
  • At Left 1 is - w.r.t. At Left 2, and this isn't mirrored in Right - Devi says this is beam related, and HVs were off...
  • Main Detectors R vs. L correlations shift in time? -> Kent says this is normal and has to do with beam motion sensitivity - HALOG
  • LHRS Dipole moved a bit - HALOG, [1]

Other Status Updates

  • Beam Mod is OK - HALOG
  • Aq is working, nothing more than ~0.2 ppm mean charge asymmetry over a given run, many sub 0.02 ppm - See Online Plots
  • SAM pedestals made SAM 8 look like a negative signal, was spurious, Devi is fine with it

July 17 2019 Swing Shift

  • WAC = Cameron Clarke
  • Slug 3 (call this slug 3.1)
  • Beam Status:
    • Day had no beam
    • Swing had 4 hours of beam, with BCM DD problems still remaining, and now the BCM widths are 128 ppm too.
  • Major issue: BCM DD noise, and now BCM noise too. Main dets corrected down to 120 ppm only.
  • Moved to 70 uA current and recalibrated gains and took new pedestal/current ramp run

Production Run List (status at Midnight, end of 7/18/2019)

Slug 4 (part 2) - 50 uA

  • Good (cut first 40k events due to excursion): 3420
  • Good: 3421
  • Good: 3422

Slug 5, 70 uA (duplicated in tomorrow's notes, special 70 uA test slug, low stats)

Changed to 70 uA!

Non-Production Runs and Bad Runs

  • 50 uA, Suspicious, don't include in run list: 3419
  • Pedestal/Current Scan: 3424 - scanned up to and around 70 uA -> Kent says its good
  • Pedestal/Current Scan: 3430 - rescanned from 10 to 70 uA

Pedestal/Calibrations Updates

Issues

Other Status Updates

  • Feedback engine gets some work to make it change more than just PITA values safely: 1, 2, 3
  • BCM DG splitter was replaced, and then unreplaced: 1
  • Bob and Silviu have changed the target encoder to properly center the target (this should not affect the blinder): HALOG